OpenStudy (anonymous):

U = Z N(x): x is a non-negative integer E(x): x is even O(x): x is odd P(x): x is prime Translate into logical notation. 1. There exists an even integer. 2. Every integer is even or odd. 3. All prime integers are non-negative. 4. The only even prime is 2. 5. Not all integers are odd. 6. Not all primes are odd. 7. If an integer is not odd, then it is even. solve any 1 or 2 parts in detail just to give me a headstart ......

OpenStudy (anonymous):

2. \[\forall U \in (E(x) \nu O(x))\] is this correct?

OpenStudy (perl):

1. $$ \Large (\exists x \in \mathbb{Z}) ~~\mathbb{E}(x) $$ if you can assume the universe is Z then $$ \Large \exists x ~\mathbb{E}(x) $$ 2. $$ \Large (\forall x \in \mathbb{Z} ) ~ ( \mathbb{E}(x) \lor \mathbb {O}(x)) $$

OpenStudy (perl):

2. if we assume U = Z integers then we have $$ \Large \forall x ~ ( \mathbb{E}(x) \lor \mathbb {O}(x)) $$

OpenStudy (anonymous):

oh kky 3. \[(\forall x \in Z ) (\forall P(x) \in N(x))\]

OpenStudy (perl):

3. $$ \Large \forall x ~( \mathbb{P}(x) \to \mathbb{N}(x)) $$

OpenStudy (perl):

where P(x) means x is prime, as you defined in your directions

OpenStudy (perl):

and you defined the universe as Z, so we can assume that is the domain of discourse

OpenStudy (anonymous):

\[what does this " \rightarrow " stands for?\]

OpenStudy (perl):

"if P then Q " $$ P \to Q $$

OpenStudy (perl):

do you use a different symbol for 'if then' ?

OpenStudy (anonymous):

no have read it in proposition logic only, didn't used it in predicate yet

OpenStudy (perl):

predicate logic includes propositional logic. predicate logic is an extension of propositional logic

OpenStudy (perl):

logical 'or' $$ \huge \lor $$ also comes from propositional logic

OpenStudy (anonymous):

\[\leftarrow \rightarrow \] will be used for only then ? in 4.

OpenStudy (perl):

the five propositional operations , 4 binary, 1 unary operator $$ \Large ~\lor ~\land~ \to ~ \leftrightarrow ~ \neg $$

OpenStudy (anonymous):

yes i can recall them now , ty

OpenStudy (anonymous):

will 4 be something like this? or am i wrong again :/ ?(\[( E(2) \leftarrow \rightarrow P(x) )\]

OpenStudy (perl):

almost

OpenStudy (perl):

4. $$ \Large \forall x ~ [(P(x) \land E(x)) \Rightarrow (x=2)] $$

OpenStudy (perl):

according to this there is a difference $$ \Large \Rightarrow \\\rightarrow $$ http://en.wikipedia.org/wiki/Material_conditional

OpenStudy (perl):

but we will use it as the simple propositional logic operator , logical conditional

OpenStudy (perl):

just plain arrow

OpenStudy (anonymous):

is it allowed to use propositional operator wit ha quantifier for e.g \[\Gamma \forall \]

OpenStudy (anonymous):

we are only allowed to use the 5 simple operators as u listed

OpenStudy (perl):

and $$ \Large \forall ~~ \exists $$ quantifiers

OpenStudy (perl):

and set membership

OpenStudy (anonymous):

yes exactly xD

OpenStudy (perl):

$$ \Large x \in y $$ is acceptable . But the way your homework is set up, we won't need to use set membership. we will need to use equal sign though. Your teacher defined E(x) as meaning $$ \Large x \in \mathbb{E} $$

OpenStudy (anonymous):

5. \[\forall x (\Gamma x \in O(x) )\]

OpenStudy (anonymous):

i hope i got it right this time .

OpenStudy (perl):

$$ \Large \mathbb{E}(x) \\ \large \text{has the same meaning as} \\ \Large x \in \mathbb{E} $$

OpenStudy (anonymous):

oh kky didn't know that

OpenStudy (perl):

5. not all integers are odd. means ' it is false that all integers are odd'

OpenStudy (perl):

or better ' it is not the case that all integers are odd' this is easiest to symbolize logically $$ \Large \neg (\forall x ~\mathbb{O}(x)) $$

OpenStudy (anonymous):

this is harder then i expected

OpenStudy (perl):

$$ \Large \neg (\forall x ~\mathbb{O}(x))\\ \Large \equiv \exists x ~\neg \mathbb{O}(x)\\ \Large \equiv \exists x ~~\mathbb{E}(x) $$

OpenStudy (anonymous):

\[\Gamma (\forall P(x) \rightarrow O(x))\]

OpenStudy (perl):

$$ \huge \equiv \\ \text{means logically equivalent, this is for the reader} $$

OpenStudy (anonymous):

it means just different ways of representing the same problem ? i can writer either one of the 3 and the answer will eb the same

OpenStudy (perl):

right,

OpenStudy (anonymous):

is the 6th one right ?

OpenStudy (perl):

yes :)

OpenStudy (perl):

your negation sign is a bit odd, capital gamma?

OpenStudy (perl):

\neg is negation

OpenStudy (anonymous):

yes xD couldn't find the negation sign so using gamma :p

OpenStudy (anonymous):

7. \[\forall x \space (\neg (O(x)) \rightarrow E(x) )\]

OpenStudy (perl):

$$ \Large \neg ~[~\forall x ~(\mathbb{P}(x)\to \mathbb{O}(x))]\\ \Large \equiv \exists x~\neg [~\mathbb{P}(x)\to \mathbb{O}(x)] \\ \Large \equiv \exists x~ (\mathbb{P}(x) \land \neg ~\mathbb{O}(x)) \\ $$

OpenStudy (perl):

here i am using the rule $$ \Large \neg (p \to q) \equiv p \land \neg~q $$

OpenStudy (perl):

useful to remember $$ \Large p \to q \equiv \neg p \lor q\\ \Large \neg (p \to q) \equiv p \land \neg~q $$

OpenStudy (anonymous):

derived from implication definition then demorgan's law application right ?

OpenStudy (perl):

right :)

OpenStudy (anonymous):

is my 7 one right ?

OpenStudy (perl):

yes just take out the parentheses around O(x)

OpenStudy (perl):

$$ \large \forall x \space (\neg O(x) \rightarrow E(x) ) $$

OpenStudy (anonymous):

oh kky thankyou very very much for your help :)

OpenStudy (perl):

this is just for reading, you did it correct

OpenStudy (perl):

good job

OpenStudy (anonymous):

thankyou :)